OpenStudy (abbysmith12):

I will medal and fan. Multiply and simplify. 1/2 x (2 1/6 + 7/8) A. 29/48 B. 1 1/24 C. 1 25/48 D. 2 25/48

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hang on :D im getting it

OpenStudy (abbysmith12):

@adrynicoleb @boricua5 @campbell_st @Directrix @eliassaab @Frostbite @Gabylovesyou @inkyvoyd @Joel_the_boss @Kainui @lizzyrockefeller @marylou004 @nechirwan

OpenStudy (anonymous):

35/16

OpenStudy (abbysmith12):

thats not a choice

OpenStudy (abbysmith12):

@undeadknight26 @ganeshie8 @willsgirl16 @Goku-Kai @dalia_lam @Notstar @IronPatriot @utkarsh3598 @Abhisar @Albany_Goon @VJT @appleduardo @loomely

OpenStudy (vjt):

what are you choices?

OpenStudy (anonymous):

yea the answer is 35x/18

OpenStudy (anonymous):

thats weird

OpenStudy (anonymous):

Step 1 : 7 7 Simplify — + — 2 8

OpenStudy (vjt):

Is this how the equation looks? \[1/2 times (2 1/6 + 7/8)\]

OpenStudy (anonymous):

Find the Least Common Multiple The left denominator is : 2 The right denominator is : 8

OpenStudy (anonymous):

Least Common Multiple: 8

OpenStudy (abbysmith12):

@abram134 @Brookie105 @c0decracker @dan815 @Ellia201 @fallenangelorchid @georgia_catherine @Harsha19111999 @iamabarbiegirl @Joel_the_boss @kevindragonlord @lizzyrockefeller @Mimi_x3 @neko_of_hearts @Ondinana @purplepiggy171 @quin100 @RAINIAR @siearra @TTP @umartouseef @VJT @Warrior678 @XceceX @yahman

OpenStudy (abbysmith12):

@ItsMeTrueAC

OpenStudy (abbysmith12):

@ilovewolf @Teddyiswatshecallsme @tester97

OpenStudy (anonymous):

yeah i'd agree with them

OpenStudy (ondinana):

I'm not so good in math and cannot help you, m apologies.. @AbbySmith12

OpenStudy (anonymous):

it has to be 35/18

OpenStudy (ondinana):

my^

OpenStudy (anonymous):

what are the choices the answer choices

OpenStudy (abbysmith12):

A. 29/48 B. 1 1/24 C. 1 25/48 D. 2 25/48

OpenStudy (abbysmith12):

@ilovewolf

OpenStudy (anonymous):

Ok its not B or D

OpenStudy (abbysmith12):

@Jamielynne

OpenStudy (anonymous):

@dan815

OpenStudy (abbysmith12):

@Radelp

OpenStudy (abbysmith12):

k

OpenStudy (anonymous):

@myronnapora

OpenStudy (abbysmith12):

@Great_And_Powerful

OpenStudy (harsha19111999):

Hey its simple. Use the BODMAS rule--Brackets Of Division Multiplication Addition Subtraction

OpenStudy (anonymous):

its C

OpenStudy (harsha19111999):

You go in order and simplify it

OpenStudy (abbysmith12):

@Darknight200

OpenStudy (anonymous):

or go to mathway.com

OpenStudy (abbysmith12):

@flasht

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

do not give direct answers

OpenStudy (harsha19111999):

You will get C. Its confirm

OpenStudy (anonymous):

use this to get your answer and you have to do the parentheses first http://www.algebrahelp.com/calculators/fractions/operations/

OpenStudy (harsha19111999):

1/2 * (13/6+7/8) = 1/2 * (73/24) = 73/48 = 1 25/48

OpenStudy (harsha19111999):

Sorry @Warrior678. I'm in a urgent work

OpenStudy (anonymous):

give*

OpenStudy (harsha19111999):

@AbbySmith12 I hope you understand

OpenStudy (anonymous):

you know who cares

OpenStudy (anonymous):

its okay @Harsha19111999

OpenStudy (anonymous):

https://www.mathway.com/ type your questions into the box and then it should help.

OpenStudy (harsha19111999):

Come on. Cool down. Warrior678 said so because AbbySmith could understand it rather than just knowing the answer

OpenStudy (anonymous):

and thats exactly why i gave you a metal

OpenStudy (anonymous):

ty abbysmith

OpenStudy (harsha19111999):

:)

OpenStudy (abbysmith12):

Whats the product? 2/3 x 5/8 x 3/4 A. 5/16 B. 5/32 C. 7/12 D. 2/3

OpenStudy (abbysmith12):

@Radelp @flasht

OpenStudy (harsha19111999):

Do you know multiplication and division?